Analysis and Interpretation of Profitability
Balance sheets and income statements for Target Corporation follow.

Income Statement
For Fiscal Years Ended ($ millions) 2007 2006 2005
Sales $ 57,878 $ 51,271 $ 45,682
Credit card revenues 1,612 1,349 1,157
Total revenues 59,490 52,620 46,839
Cost of sales 39,399 34,927 31,445
Selling, general and administrative expenses 12,819 11,185 9,797
Credit card expenses 707 776 737
Depreciation and amortization 1,496 1,409 1,259
Earnings before interest and income taxes 5,069 4,323 3,601
Net interest expense 572 463 570
Earnings before income taxes 4,497 3,860 3,031
Provisions for income taxes 1,710 1,452 1,146
Net earnings $ 2,787 $ 2,408 $ 1,885
Balance Sheet
($ millions, except footnotes) February 3, 2007 January 28, 2006
Assets
Cash and cash equivalents $ 813 $ 1,648
Credit card receivables 6,194 5,666
Inventory 6,254 5,838
Other current assets 1,445 1,253
Total current assets 14,706 14,405
Property and equipment
Land 4,934 4,449
Buildings and improvements 16,110 14,174
Fixtures and equipment 3,553 3,219
Computer hardware and software 2,188 2,214
Construction-in-progress 1,596 1,158
Accumulated depreciation (6,950) (6,176)
Property and equipment, net 21,431 19,038
Other noncurrent assets 1,212 1,552
Total assets $ 37,349 $ 34,995
Liabilities and shareholders' investment
Accounts payable $ 6,575 $ 6,268
Accrued and other current liabilities 3,180 2,567
Current portion of long-term debt and notes payable 1,362 753
Total current liabilities 11,117 9,588
Long-term debt 8,675 9,119
Deferred income taxes 577 851
Other noncurrent liabilities 1,347 1,232
Shareholders' investment
Common stock 72 73
Additional paid-in-capital 2,387 2,121
Retained earnings 13,417 12,013
Accumulated other comprehensive income (loss) (243) (2)
Total shareholders' investment 15,633 14,205
Total liabilities and shareholders' equity $ 37,349 $ 34,995


a. Compute ROE for 2007.

Do not round until your final answer. Round answer to two decimal places.

ROE =Answer%


b. Confirm that ROE equals ROE computed using the component measures for profit margin, asset turnover, and financial leverage using: ROE = PM * AT * FL.

   Compute the components of ROE.

   Do not round until your final answer. Round answers to two decimal places.

   PM = Answer%
   AT = Answer
   FL = Answer

  
c. Compute adjusted ROA. Assume a tax rate of: 39.0%.

Do not round until your final answer. Round your answer to two decimal places.
Adjusted ROA =Answer%
Answer #1
Part a
ROE = Net earnings / Shareholders equity
ROE = 2787 / 15633
ROE = 17.83%
Part b
Profit margin = 2787 / 57878 = 4.82%
Asset turnover = 57878 / 37349 = 1.55
Finacial leverage = 37349 / 15633 = 2.39
ROE = PM x AT x FL
ROE = 4.82% * 1.55 * 2.39
ROE = 17.83%
Part c
Adjusted ROA = Adjusted Earnings / Adjusted Total assets
Adjusted ROA = (4497*(1-39%)) / 37349
ROA = 7.34%